Sol Summary – April 19th

Crew 245 Sol Summary Report 19-04-2022
SOL: 9
Summary Title: Nominal ops
Author’s name: Paolo Guardabasso

Mission status: The majority of the experiments proceed nominally, with some issues here and there.

Sol Activity Summary: The day was dedicated to EVAs in the morning (before the wind rose) and to sample and data processing in the afternoon, with a bit of relax. We also (sadly) started managing the logistics for our return to Earth.

Look Ahead Plan: Tomorrow we’ll once again dedicate the morning to EVAs, and spend the afternoon most likely similar to today, working on experiments and logistics.

Anomalies in work: We observed a non-nominal behaviour in Suit 1 (air only at maximum), and we performed some troubleshooting.

Weather: Good weather, although increasingly windy. Temperature 9 (low) and 29 (high) C, Wind speed 43 (avg) and 64 (max) kph

Crew Physical Status: We’re starting to feel the effects of the hot season. One crew member was denied EVA because of high BP.

EVA: Two consecutive EVAs have been performed. They represent the first part of two parts of the stress level experiment, where 2 crew members will have to explore and navigate an area with and without a digital map previously generated with a drone.

Reports to be filled:
– HSO report
– Operation report
– Journalist report
– Some pictures + picture of the day
– EVA request (2)
– EVA report (2)

Support Requested: None at the moment.

Copyright © The Mars Society. All rights reserved. | Main Site